OpenStudy (anonymous):

determine the value of 'x' that makes each sequence arithmetic. d) x-4, 6, x e) x+8, 2x+8, -x

OpenStudy (turingtest):

if each term in the sequence is equidistant from the next we can write the difference between each term mathematically and set them equal\[6-(x-4)=x-6\]solve for x

OpenStudy (turingtest):

...because each term is equidistant from the next in an arithmetic progression

OpenStudy (anonymous):

so then 6-x+4=x-6 10-x=x-6 10+6=x+x 16=2x 8=x yay ! thanks :)

OpenStudy (turingtest):

welcome :D same deal for the next one as well

OpenStudy (anonymous):

so it would be 2x+8 - (x+8) = -x right ?

OpenStudy (turingtest):

you forgot the difference of the last terms 2x+8 - (x+8) = -x - (2x +8)

OpenStudy (anonymous):

nvm,i got it,thanks man! its -2

OpenStudy (turingtest):

call the entries of the series\[E_1,E_2,E_3\]if it is an arithmetic progression we have\[E_2-E_1=E_3-E_2\]is the idea... again, welcome!